The box having mass “m” is being pulled by a motor.Initial velocity of the box is V0.Motor pulles the box by force F=500s+550,where “s” is the distance the box is pulled (meters),and force is in newtons.Define the velocity and acceleration of the box as a function of distance ‘s’.If the motor has efficiency of “E(epsilon)=0,65” ,power input required to be supplied by the motor as a function of distance.
1
Expert's answer
2019-11-20T09:54:58-0500

According to Newton’s second law, acceleration "a = \frac{500s+550}{m},"


and since "m(\frac{dv}{ds})(\frac{ds}{dt}) = F," then "V = (\frac{500s^2 + 1100s}{m} + v_0^2)^{0.5}."